OpenStudy (anonymous):

Medal and fan :3 Find all the real square roots of -9/16

OpenStudy (mathstudent55):

-9/16 is a negative number. What is 2 * 2 = ? What is -2 * (-2) = ?

OpenStudy (tkhunny):

\(x^{2} = -\dfrac{9}{16}\) \(x^{2} + \dfrac{9}{16}= 0\) No positive Real Roots. No negative Real Roots.

OpenStudy (mathmale):

If there are neither positive real roots nor negative real roots, what have you (@Justmeneedinghelp )?